OpenStudy (anonymous):

Solve for X 0=5x^2 +10x -120

OpenStudy (callisto):

0=5x^2 +10x -120 0=5(x^2+2x-24) 0=x^2+2x-24 0=(x+6)(x-4) (x+6) = 0 or (x-4)=0 x = -6 or x=4